# TextCenGen: Attention-Guided Text-Centric Background Adaptation for Text-to-Image Generation

### v3

@@ Abstract
Recent advancements in Text-to-image (T2I) generation have witnessed a shift from adapting text to fixed backgrounds to creating images around text. Traditional approaches are often limited to generate layouts within static images for effective text placement. Our proposed approach, TextCenGen, introduces a dynamic adaptation of the space region for text-friendly image generation, focusing on text-centric design and visual harmony generation. Our method employs force-directed attention guidance in T2I models to generate images that strategically reserve whitespace for pre-defined text areas, even for text or icons at the golden ratio. Observing how cross-attention maps affect object placement, we detect and repel conflicting objects using a force-directed graph approach, combined with a spatial excluding cross-attention constraint for smooth attention in whitespace areas. As a novel task in graphic design, experiments indicate that TextCenGen outperforms existing methods with more harmonious compositions. Furthermore, our method significantly enhances T2I model outcomes on our specially collected prompt datasets, catering to different text positions. These results demonstrate the efficacy of TextCenGen in creating more harmonious and integrated text-image compositions.
Figure 1: TextCenGen is a training-free method designed to generate text-friendly images. It simultaneously generates the original and result images. During the denoising process of the original image, the generation of the result image is guided based on the encroachment into the planned space regions in the original. This approach ensures sufficient blank space at specific positions, typically where text or icons are centered, in the result image.
Figure 2: Our target text-friendly images. (a) An input layout for text at the golden ratio, intended as a Valentine’s Day card on a selected background image, (b) shows background interference causing unappealing text visibility, and (c) presents standard text-friendly image adaptation with clear text integration.
1 Introduction
@@
In graphic design, achieving an effective synergy between text and imagery is essential for clear communication. The choice of background can greatly influence text visibility and comprehension. A common design objective is to place text within an image in a way that is both visually pleasing and communicatively clear. A preferred strategy is to position the text in the golden ratio, which is believed to be aesthetically optimal (see Figure 2a). However, designers often grapple with the issue of backgrounds that compete with or obscure the text, as illustrated in Figure 2b, where unsuitable backgrounds detract from the text’s readability and aesthetic appeal, regardless of any adjustments to text color or size. We aim to facilitate the creation of text-friendly images (see Figure 2c), ideal for text placement and catering to the growing demand due to increasing use of Text-to-Image (T2I) models for background graphics.
@@
Traditional approaches to graphic design, especially in poster creation, have largely focused on arranging layouts with natural background images, elements, and text Guo et al. (2021); Cao et al. (2012); O’Donovan et al. (2014); Li et al. (2022). However, producing text-friendly images remains a significant challenge due to the complexity of background elements. Our insight, derived from Figure 2c, reveals that a clear separation between the main objects and the text areas is essential. Recent advancements in diffusion-related research have shown that it is possible to manipulate the primary objects within cross-attention maps, making the adaptation of background images to accommodate text a feasible endeavorHertz et al. (2022); Epstein et al. (2023).
@@
In response, we introduce TextCenGen, as illustrated in Figure 1, a novel method that employs cross-attention maps and force-directed graphs for effective object placement and whitespace optimization. We also implement a spatial excluding cross-attention constraint to ensure smooth attention in areas designated for text. To establish a new benchmark for this innovative task, we constructed a diverse dataset gathered from three unique sources, along with three distinct evaluation metrics, to comprehensively assess the performance. The contributions of our paper are three-fold:
@@
  • We propose a novel task of text-friendly text-to-image generation, which creates images that satisfy both the prompt and reserve space for pre-defined text placements. Our benchmark includes a specialized dataset and tailored evaluation metrics designed to rigorously assess this novel challenge.
@@ 2 Related Work
Text Layout of Natural Images has evolved significantly, transitioning from traditional layout designs to more advanced methods influenced by deep learning. Initially, poster design focused on creating layouts with given background images, elements, and text Guo et al. (2021); Cao et al. (2012); O’Donovan et al. (2014); Li et al. (2022). The integration of deep learning in text layout of natural image generation has led to various models such as GAN Goodfellow et al. (2014); Zheng et al. (2019); Li et al. (2019); Zhou et al. (2022), VAE Jyothi et al. (2019), transformers Vaswani et al. (2017); Inoue et al. (2023b); Wang et al. (2023) and diffusion models Ho et al. (2020); Hui et al. (2023); Inoue et al. (2023a); Chai et al. (2023); Li et al. (2023a). These models have been instrumental in learning layout patterns from large datasets. Subsequent research explored methods to retrieve matching background images based on text and image description Jin et al. (2022). Since the development of image editing method based on diffusion model, scene text generation methods such as TextDiffuser Chen et al. (2023b, a) and DiffText Zhang et al. (2023a) have addressed the challenges of generating clear text with diffusion models. However, these methods often rely on the presence of a ’sign’ or similar element within the prompt (e.g., a T-shirt) to place text. They do not explicitly tackle the problem of creating text-friendly images where the background itself is crafted to adapt pre-defined text regions. Our approach, TextCenGen, significantly extends these capabilities by allowing the primary objects in generated images to yield space to text regions, resulting in more harmonious and aesthetically pleasing compositions.
@@
Text-to-Image Generation has advanced with diffusion models Ho et al. (2020), producing realistic images and videos that align closely with text prompts Rombach et al. (2022); Singer et al. (2023); Chen et al. (2023c). Innovations in this field include GLIDE Nichol et al. (2022), which integrates text conditions into the diffusion process, DALL-E 2 Ramesh et al. (2022) with its diffusion prior module for high-resolution images, and Imagen Saharia et al. (2022), which uses a large T5 language model to enhance semantic representation. Stable diffusion Rombach et al. (2022) projects images into latent space for diffusion processing. Beyond text conditions, the manipulation of diffusion models through image-level conditions has been explored. Methods such as image inpainting Balaji et al. (2022) aim to generate coherent parts of an image, while SDG Liu et al. (2023) introduces semantic inputs to guide unconditional DDPM sampling. In addition, techniques such as Meng et al. (2021) use images as editing conditions in denoising processes. Scene text generation methods such as TextDiffuser Chen et al. (2023b, a) and GlyphDraw Ma et al. (2023) have also emerged, utilizing textual layouts or masks to guide text generation in images. These advancements represent the growing versatility and potential of T2I models in diverse applications.
@@
Attention Guided Image Editing has emerged as a critical solution to the challenge of translating human preferences and intentions into visual content through text descriptions. These approaches, as seen in works like Li et al. (2023c); Avrahami et al. (2023); Zhang et al. (2023b); Zhao et al. (2023), involve learning auxiliary modules on paired data. However, a significant limitation of these training-based methods is the substantial cost and effort required for repeated training for different control signals, model architectures, and checkpoints. In response to these challenges, training-free techniques have emerged, using the inherent weights of attention and the characteristics of pre-trained models to control the characteristics of objects such as size, shape, appearance, and location Hertz et al. (2022); Epstein et al. (2023); Patashnik et al. (2023); Xie et al. (2023). These methods typically utilize basic conditions, such as bounding boxes, for precise control over object positioning and scene composition Mo et al. (2023). Our approach takes this further by applying force-directed graph techniques to cross-attention map edits, allowing for more automated and precise object transformations in T2I editing.
3 Method
@@ 3.1 Overview
Initially, we consider the $k^{th}$ token of $T_{prompt}$, which corresponds to the cross-attention map $A_{k}$. Some methods utilize $A_{k}$ to locate and edit objects within generated images Epstein et al. (2023). In response to our first concern, we analyze the denoising process of an unguided image ($I_{ori}$) from a given prompt. We establish the subset of tokens ($O$) that refer to objects that need modification. The conflict detector operates based on the average attention intensity shared between the shapes of objects corresponding to $A_{k}$ and $R$ in an overlapping manner. For every token $k\in O$, we introduce Force-Directed Cross-Attention Guidance for moving objects. In this scheme, objects are treated as centroid vertex within a graph, with a sequence of forces being applied (see  Figure 4) to adjust object positions.
For the second consideration, we employ strategies derived from recent technologies that limit the range of the attention map Zhang et al. (2023a). Consequently, we propose a spatial excluding cross-attention constraint to prevent an extensive attention density from encroaching on $R$.
We find that repulsive force often moves required objects out of the image. Addressing the third concern, our denoising process incorporates a loss function with additional regularization terms to safeguard the shapes and positions of other objects. To prevent objects from being dislocated outside limits while retaining their reasonable shapes, we also introduce the notions of Margin Force $F_{m}()$ and Warping Force $F_{w}()$.
3.2 Force-Directed Cross-Attention Guidance
Figure 4: Illustration of four set relationships and their associated forces. The Repulsive Force separates object and text centroids during intersections (a1) and enclosures (a2). The Margin Force (b) and Warping Force (c) prevent boundary overstepping. Details on text exclusion within object regions (a4) are discussed in the ’Excluding Cross-Attention Constraint’ section.
@@ 3.2 Force-Directed Cross-Attention Guidance
To seamlessly integrate the concept of force-directed graphs into the loss guidance of the denoising process in latent diffusion models, we delve into the extraction and manipulation of attention maps and activations. We use softmax normalized attention matrices $\mathcal{A}_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times K}$ extracted from the standard denoising forward step $\epsilon_{\theta_{i}}(z_{i};t,y)$. This enables us to manipulate the control over objects referred to in the text conditioning $y$ at distinct indices $k$, by adjusting the related attention channel(s) $\mathcal{A}_{i,t,\ldots,k}\in\mathbb{R}^{H_{i}\times W_{i}\times|k|}$. The centroid of the attention map is a two-dimensional vector, defined by the equation: 1
@@
where $\omega_{j}$ are coefficients for balancing attention across targets. To regulate the impact of these forces and avoid excessive dominance by any single target, we introduce a force balance constant $\alpha$ in the form of $\frac{F_{rep}}{\alpha+F_{rep}}$. $\alpha$ ensures that the forces exerted do not exceed a practical threshold, thereby maintaining visual equilibrium in complex scenes (shown in  Figure 5).
Figure 5: Influence of the force balance constant. The graph illustrates the effects of four different values for $\alpha$ on $F_{rep}$ versus the final force. The curves reveal that as $\alpha$ increases, the rate of convergence to the limit progresses more rapidly.
@@
To compute the total displacement $\Delta pos(v)$, we sum the repulsive and margin forces together: $\Delta pos(v)=F_{rep}(v)+F_{m}(v)$ Subsequently, we update the vertex’s position as follows: $pos(v)=pos(v)+\Delta pos(v)$. For the attention map $A_{k}^{l}$, this update is applied as a whole, with excess regions outside the boundaries being discarded and the remaining areas filled with zeros. But this method introduces the risk of the object being moved out of the boundaries and then being discarded, so it is necessary to introduce the following ’warping force’ to prevent this from happening.
@@
In addressing the dynamics of our force-directed graph algorithm, particularly for the movement of cross-attention maps $\mathcal{A}_{k}^{l}$ in each layer, we employ affine transformations as a pivotal mechanism. This approach facilitates the comprehensive translation and scaling of the entire map, preserving the relative positions within the image domain. Initially, we delineate our visual area or image space as a $H\times W$ two-dimensional array $A$. Within this canvas $A$, we identify a key region, the object $O$, defined by the coordinates $(x,y,a,b)$, where $(x,y)$ marks the upper-left corner and $(a,b)$ the lower-right corner. The movement (move) is calculated based on the sum of repulsive force $F_{\text{rep}}(v)$ and margin force $F_{\text{m}}(v)$, yielding the total displacement $\Delta\text{pos}(v)=F_{\text{rep}}(v)+F_{\text{m}}(v)$. Applying $\Delta\text{pos}(v)$ to both $A$ and $O$, we obtain the transformed canvas $A^{\prime}$ and the object $O^{\prime}$. Subsequently, we shift our coordinate system’s origin to a vertex within $A^{\prime}$ that remains within the canvas boundaries, establishing a new origin $O_{\text{new}}$. This repositioning is crucial when $O^{\prime}$ exceeds the visual boundaries of $A$. In such cases, we scale the moved $\mathcal{A}_{k}^{l}$ to ensure that the bounding box of $O^{\prime}$ fits precisely within the confines of $\mathcal{A}_{k}^{l}$. Scaling factors $S_{x}$ and $S_{y}$, are calculated as $S_{x}=\min\left(1,\frac{H-1}{a^{\prime}}\right)$ and $S_{y}=\min\left(1,\frac{W-1}{b^{\prime}}\right)$, where $(a^{\prime},b^{\prime})$ are the new coordinates of $O^{\prime}$. Finally, the scaled $\mathcal{A}_{k}^{l}$ and $O^{\prime}$ are reverted back to their original coordinate origin (see the warping force in  Figure 4). A critical aspect of our approach is the transformation of the reference frame. After displacing $O$ due to $\Delta\text{pos}(v)$, a new reference frame is established, centered at $O_{\text{new}}$. The coordinates of $O$ in this new frame are calculated as $(x_{\text{new}},y_{\text{new}})=(x+\Delta x-O_{\text{new}_{x}},y+\Delta y-O_{\text{new}_{y}})$. The affine transformation, considering this frame shift, is represented as:
@@
3.3 Spatial Excluding Cross-Attention Constraint
Our goal is to maintain a smooth background in the text region denoted as $R$. As illustrated in Figure 3, during each timestep of the forward pass in the diffusion model, we modify the cross-attention maps at every layer. The cross-attention map is represented as $\mathcal{A}_{k}^{l}\in\mathbb{R}^{H\times W\times K}$, where $H\times W$ are the dimensions of $A^{l}$ at different scales, and $K$ signifies the maximum token length at layer $l$. The set $I$ consists of indices of tokens corresponding to areas outside the text in the prompt. We resize $R$ to align with the $HW$ dimensions. Subsequently, a new cross-attention map for each layer $l$ is computed as $\mathcal{A}_{k,\text{new}}^{l}=\{\lambda\times\mathcal{A}_{k}^{l}\odot(1-R)\,|\,\forall k\in O\}$, where $O$ represents the tokens needing editing. This procedure effectively redirects the model’s attention away from $R$, ensuring that the background in this region remains undisturbed and visually smooth. This spatially exclusive approach enhances the clarity and coherence of the generated images, particularly in areas designated for text insertion.
The final loss function of our denoising process is defined as:
@@ 3.3 Spatial Excluding Cross-Attention Constraint
In this case, $\mathcal{A}_{ori}$ and $\mathcal{A}_{res}$ stand for the original and altered versions of matrix $A$, correspondingly, while ||.||||.||| | . | | signifies the Euclidean distance among these cross-attention maps. Furthermore, the inclusion of the regularization term $\gamma\mathcal{L}_{norm}$ derived from self-guidanceEpstein et al. (2023), strives to prevent overfitting, where $\gamma$ denotes a delicately selected weighting parameter intended to maintain a balance in the impact of the norm penalty, $\mathcal{L}_{norm}$, within the aggregate loss function $\mathcal{L}$.
Figure 6: The results of comparison. Each column showcases six varied prompts, the final column depicting the saliency map of the result image generated from the mushroom prompt. The red-dotted area denotes the target area preserved for either text or icon images.
Figure 7: The results of ablation. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The red-dotted area denotes the target area preserved for either text or icon images. Images produced with both FDG and SEC yield outcomes identical to those created using TextCenGen alone. Conversely, images created without FDG and SEC are comparable to those derived from the original stable diffusion model.
 Dataset Metrics Stable Diffusion Dall-E 3 AnyText DiffText Ours
P2P Template CLIP Score $\uparrow$ 28.26 25.73 27.94 17.67 27.96
Saliency IOU $\downarrow$ 29.89 52.64 54.34 21.30 28.37
TV Loss $\downarrow$ 14.11 18.02 22.55 9.10 13.29
VTCM $\uparrow$ 2.95 1.92 1.75 2.77 3.09
DiffuisonDB CLIP Score $\uparrow$ 27.84 25.80 26.80 17.92 27.20
Saliency IOU $\downarrow$ 30.11 56.82 55.88 23.49 27.80
TV Loss $\downarrow$ 12.30 21.16 20.39 8.86 11.55
VTCM $\uparrow$ 2.73 1.98 1.79 2.41 2.76
Synthetics Prompt CLIP Score $\uparrow$ 28.45 26.20 27.43 18.78 28.10
Saliency IOU $\downarrow$ 31.42 51.52 53.24 24.00 30.83
TV Loss $\downarrow$ 15.61 17.85 21.46 11.56 15.17
VTCM $\uparrow$ 3.19 1.67 1.79 2.79 3.33
Table 1: Quantitative comparison of our method with other methods. Red indicates the best scores and blue marking the second best scores.
4 Experiments
Our evaluation strategy is designed to thoroughly assess the effectiveness of TextCenGen in generating text-centric images. The evaluation is structured into quantitative and qualitative analyses, alongside an ablation study to understand the contribution of individual components of our model.
4.1 Implementation Details
@@
Dataset for Evaluation
Our evaluation contains 27,000 images generated from 2,700 unique prompts, each tested in ten different random region $R$. The dataset combined synthesized prompts generated by ChatGPT, and 700 prompts from the Prompt2Prompt template ( Hertz et al. (2022)) designed for attention guidance, focusing on specific objects and their spatial relationships. Additionally, we included 1,000 DiffusionDB prompts ( Wang et al. (2022)), chosen for their real-world complexity. This diverse and comprehensive dataset, spanning synthetic to user-generated prompts, provided a broad test ground to evaluate the efficacy of our TextCenGen method in various T2I scenarios.
4.2 Comparison with Existing Methods
We compared TextCenGen with several potential models to evaluate its efficiency. The baseline models include: Native Stable Diffusion Rombach et al. (2022), Dall-E 3 Ramesh et al. (2022), AnyText Tuo et al. (2023) and DiffText Zhang et al. (2023a). Dall-E uses the prompt ”text-friendly in the position” to specify the region $R$. Similar to AnyText and DiffText, we chose to randomly generate several masks in a fixed pattern across the map to simulate regions need to be edited. More detail can be found in appendix.

### v4

@@ Abstract
Recent advancements in Text-to-Image (T2I) generation have witnessed a shift from adapting text to fixed backgrounds to creating images around text. Traditional approaches are often limited to generate layouts within static images for effective text placement. Our proposed approach, TextCenGen, introduces a training-free and plug-and-play dynamic background adaptation of the blank region for text-friendly image generation. Our method employs force-directed attention guidance in T2I models to generate images that strategically reserve whitespace for any pre-defined text areas. Observing how cross-attention maps affect object placement, we detect and repel conflicting objects using a force-directed graph approach, combined with a spatial excluding cross-attention constraint for smooth attention in whitespace areas. To evaluate our method, we propose hybrid prompt datasets as benchmarks with . Experiments indicate that TextCenGen outperforms existing methods by producing more harmonious compositions, demonstrating that it effectively enhances T2I model outcomes by catering to different text positions.
Figure 1: TextCenGen is a training-free method designed to generate text-friendly images. By using a simple text prompt and a planned blank region as inputs, TextCenGen creates images that satisfy the prompt and provide sufficient blank space in the target region. For example, the text-friendly T2I approach helps users customize their favored text-friendly wallpapers for mobile devices with any T2I model, avoiding visual confusion caused by main objects overlapping with UI components.
Introduction
111*These authors contributed equally to this work222†Corresponding author: chli@cs.ecnu.edu.cn
@@ Introduction
In graphic design, achieving an effective synergy between text and imagery is essential for clear communication. The choice of background can influence text visibility and comprehension. A common design objective is to place text within an image in a way that is both visually pleasing and communicatively clear. A preferred strategy is to position the text in the golden ratio, which is believed to be aesthetically optimal. However, designers often grapple with the issue of backgrounds that compete with or obscure the text, as illustrated in Figure 1[Uncaptioned image], where unsuitable backgrounds detract from the text’s readability and aesthetic appeal, regardless of any adjustments to text color or size. We aim to facilitate the creation of text-friendly images (see Figure 1[Uncaptioned image]), ideal for text placement and meeting the growing demand due to increasing use of Text-to-Image (T2I) models for background graphics.
@@
Traditional approaches to graphic design, especially in poster creation, have largely focused on arranging layouts in static natural background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). However, producing text-friendly images remains a significant challenge due to the complexity of background elements. Our insight, derived from Figure 1, reveals that a clear separation between the main objects and the text areas is essential. Recent advancements in diffusion-related research have shown that it is possible to manipulate the primary objects within cross-attention maps, making the adaptation of background images to accommodate text a feasible endeavor (Hertz et al. 2022; Epstein et al. 2023).
@@
In response, we introduce TextCenGen[Uncaptioned image], as illustrated in Figure 1, a novel method that employs cross-attention maps and force-directed graphs for effective object placement and whitespace optimization. We also implement a spatial excluding cross-attention constraint to ensure smooth attention in areas designated for text. To establish a new benchmark for this innovative task, we constructed a diverse dataset gathered from three unique sources, along with three distinct evaluation metrics, to comprehensively assess the performance. The contributions of our paper are three-fold:
@@
  • We propose a new task of text-freindly T2I generation, which creates images that satisfy both the prompt and reserve space for pre-defined text placements. The task consists of a benchmark including a specialized dataset and tailored evaluation metrics.
@@
Text Layout of Natural Images has evolved significantly, transitioning from traditional layout designs to more advanced methods influenced by deep learning. Initially, poster design focused on creating layouts with given background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). The integration of deep learning in text layout of natural image generation has led to various models such as GAN (Goodfellow et al. 2014; Zheng et al. 2019; Li et al. 2019; Zhou et al. 2022), VAE (Jyothi et al. 2019), transformers (Vaswani et al. 2017; Inoue et al. 2023b; Wang et al. 2023) and diffusion models (Ho, Jain, and Abbeel 2020; Hui et al. 2023; Inoue et al. 2023a; Chai, Zhuang, and Yan 2023; Li et al. 2023a). These models have been instrumental in learning layout patterns from large datasets. Subsequent research explored methods to retrieve matching background images based on text and image description (Jin et al. 2022). Since the development of image editing method based on diffusion model, scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and DiffText (Zhang et al. 2023) have addressed the challenges of generating clear text with diffusion models. However, these methods often rely on the presence of a “sign” or similar element within the prompt (e.g., a T-shirt) to place text. They do not explicitly tackle the problem of creating text-friendly images where the background itself is crafted to adapt pre-defined text regions. Our approach extends these capabilities by allowing the primary objects in generated images to yield space to text regions, resulting in more harmonious and aesthetically pleasing compositions.
@@
Text-to-Image Generation has advanced with diffusion models (Ho, Jain, and Abbeel 2020), producing realistic images and videos that align closely with text prompts (Rombach et al. 2022; Singer et al. 2023; Chen et al. 2023c). Innovations in this field include GLIDE (Nichol et al. 2022), which integrates text conditions into the diffusion process, Dall-E 2 (Ramesh et al. 2022) with its diffusion prior module for high-resolution images, and Imagen (Saharia et al. 2022), which uses a large T5 language model to enhance semantic representation. Stable diffusion (Rombach et al. 2022) projects images into latent space for diffusion processing. Beyond text conditions, the manipulation of diffusion models through image-level conditions has been explored. Methods such as image inpainting (Balaji et al. 2022) aim to generate coherent parts of an image, while SDG (Liu et al. 2023) introduces semantic inputs to guide unconditional DDPM sampling. In addition, techniques such as (Meng et al. 2021) use images as editing conditions in denoising processes. Scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and GlyphDraw (Ma et al. 2023) have also emerged, utilizing textual layouts or masks to guide text generation in images. These advancements represent the growing versatility and potential of T2I models in diverse applications.
@@
Attention Guided Image Editing has emerged as a foundamental solution to the challenge of translating human preferences and intentions into visual content through text descriptions. These approaches, as seen in works like (Li et al. 2023b; Avrahami et al. 2023; Zhang, Rao, and Agrawala 2023; Zhao et al. 2023), involve learning auxiliary modules on paired data. However, a significant limitation of these training-based methods is the substantial cost and effort required for repeated training for different control signals, model architectures, and checkpoints. In response to these challenges, training-free techniques have emerged, using the inherent weights of attention and the characteristics of pre-trained models to control the characteristics of objects such as size, shape, appearance, and location (Hertz et al. 2022; Epstein et al. 2023; Patashnik et al. 2023; Xie et al. 2023). These methods typically utilize basic conditions, such as bounding boxes, for precise control over object positioning and scene composition (Mo et al. 2023). Desigen (Weng et al. 2024) discovered relationship between attention and saliency and introduce attention reduction to weaken the attention whin layout boxes. Our approach takes this further by applying force-directed graph techniques to cross-attention map edits, allowing for more automated and precise object transformations in T2I editing.
Method
@@
As cross-attention map $A_{k}$ serves as a medium to locate and edit objects within generated images (Epstein et al. 2023), we focus on the $k^{th}$ token of $T_{prompt}$. In response to our first concern, we analyze the denoising process of an unguided image ($I_{ori}$) from a given prompt, and establish the subset of tokens ($O$) that refer to objects that need modification. We design a conflict detector that determines object conflicts based on the average attention intensity in the overlapping regions between $A_{k}$ and $R$. For every token $k\in O$, we introduce Force-Directed Cross-Attention Guidance for moving objects. In this scheme, objects are treated as centroid vertex within a graph, with a sequence of forces being applied (see Figure 3) to adjust object positions.
For the second consideration, inspired by recent technologies that limit the range of the attention map (Zhang et al. 2023), we propose a spatial excluding cross-attention constraint to prevent an extensive attention density from encroaching on $R$.
Addressing the third concern, our denoising process incorporates a loss function with additional regularization terms to safeguard the shapes and positions of other objects. Sometimes excessive repulsive force can occasionally displace essential objects from the image. To prevent objects from being dislocated outside limits while retaining their reasonable shapes, we also introduce the notions of Margin Force $F_{m}()$ and Warping Force $F_{w}()$.
Force-Directed Cross-Attention Guidance
Figure 3: Illustration of four set relationships and their associated forces. The Repulsive Force separates object and text centroids during intersections (a1) and enclosures (a2). The Margin Force (b) and Warping Force (c) prevent boundary overstepping. Text within object regions (a4) requires cooperation between force and attention constraint.
@@
To seamlessly integrate the concept of force-directed graphs into the loss guidance of the denoising process in latent diffusion models, we delve into the extraction and manipulation of attention maps and activations. We use softmax normalized attention matrices $\mathcal{A}_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times K}$ of layer $i$ extracted from the standard denoising forward step $\epsilon_{\theta_{i}}(z_{i};t,y)$. This enables us to manipulate the control over objects referred to in the text conditioning $y$ at distinct indices $k$, by adjusting the related attention channel(s) $\mathcal{A}_{i,t,\ldots,k}\in\mathbb{R}^{H_{i}\times W_{i}\times|k|}$. The centroid of the attention map is a two-dimensional vector, defined by the equation:
@@
where $\omega_{j}$ are coefficients for balancing attention across targets. To regulate the impact of these forces and avoid excessive dominance by any single target, we introduce a force balance constant $\alpha$ in the form of $\frac{F_{rep}}{\alpha+F_{rep}}$. $\alpha$ ensures that the forces exerted do not exceed a practical threshold, thereby maintaining visual equilibrium in complex scenes.
@@
To compute the total displacement $\Delta pos(v)$, sum the repulsive and margin forces together: $\Delta pos(v)=F_{rep}(v)+F_{m}(v)$ Subsequently, update the vertex’s position as follows: $pos(v)=pos(v)+\Delta pos(v)$. For the attention map $A_{k}^{l}$, this update is applied as a whole, with excess regions outside the boundaries being discarded and the remaining areas filled with zeros. But this method introduces the risk of the object being moved out of the boundaries and then being discarded, so it is necessary to introduce the following ’warping force’ to prevent this from happening.
@@
In addressing the dynamics of our force-directed graph algorithm, particularly for the movement of cross-attention maps $\mathcal{A}_{k}^{l}$ in each layer, we employ affine transformations as a pivotal mechanism. This approach facilitates the comprehensive translation and scaling of the entire map, preserving the relative positions within the image domain. Initially, we delineate our visual area or image space as a $H\times W$ two-dimensional array $A$. Within this canvas $A$, we identify a key region, the object $O$, defined by the coordinates $(x,y,a,b)$, where $(x,y)$ marks the upper-left corner and $(a,b)$ the lower-right corner. The movement is calculated based on the sum of repulsive force $F_{\text{rep}}(v)$ and margin force $F_{\text{m}}(v)$, yielding the total displacement $\Delta\text{pos}(v)=F_{\text{rep}}(v)+F_{\text{m}}(v)$. Applying $\Delta\text{pos}(v)$ to both $A$ and $O$, we obtain the transformed canvas $A^{\prime}$ and the object $O^{\prime}$. Subsequently, we shift our coordinate system’s origin to a vertex within $A^{\prime}$ that remains within the canvas boundaries, establishing a new origin $O_{\text{new}}$. This repositioning is crucial when $O^{\prime}$ exceeds the visual boundaries of $A$. In such cases, we scale the moved $\mathcal{A}_{k}^{l}$ to ensure that the bounding box of $O^{\prime}$ fits precisely within the confines of $\mathcal{A}_{k}^{l}$. Scaling factors $S_{x}$ and $S_{y}$, are calculated as $S_{x}=\min\left(1,\frac{H-1}{a^{\prime}}\right)$ and $S_{y}=\min\left(1,\frac{W-1}{b^{\prime}}\right)$, where $(a^{\prime},b^{\prime})$ are the new coordinates of $O^{\prime}$. Finally, the scaled $\mathcal{A}_{k}^{l}$ and $O^{\prime}$ are reverted back to their original coordinate origin (see the warping force in Figure 3). A critical aspect of our approach is the transformation of the reference frame. After displacing $O$ due to $\Delta\text{pos}(v)$, a new reference frame is established, centered at $O_{\text{new}}$. The coordinates of $O$ in this new frame are calculated as $(x_{\text{new}},y_{\text{new}})=(x+\Delta x-O_{\text{new}_{x}},y+\Delta y-O_{\text{new}_{y}})$. The affine transformation, considering this frame shift, is represented as:
@@
Spatial Excluding Cross-Attention Constraint
Our goal is to maintain a smooth background in the text region denoted as $R$. As illustrated in Figure 2, during each time-step of the forward pass in the diffusion model, we modify the cross-attention maps at every layer. The cross-attention map is represented as $\mathcal{A}_{k}^{l}\in\mathbb{R}^{H\times W\times K}$, where $H\times W$ are the dimensions of $A^{l}$ at different scales, and $K$ signifies the maximum token length at layer $l$. The set $I$ consists of indices of tokens corresponding to areas outside the text in the prompt. We resize $R$ to align with the $HW$ dimensions. Subsequently, a new cross-attention map for each layer $l$ is computed as $\mathcal{A}_{k,\text{new}}^{l}=\{\lambda\times\mathcal{A}_{k}^{l}\odot(1-R)\,|\,\forall k\in O\}$, where $O$ represents the tokens needing editing. This procedure effectively redirects the model’s attention away from $R$, ensuring that the background in this region remains undisturbed and visually smooth. This spatially exclusive approach enhances the clarity and coherence of the generated images, particularly in areas designated for text insertion.
The final loss function of denoising process is defined as:
@@
In this case, $\mathcal{A}_{ori}$ and $\mathcal{A}_{res}$ stand for the original and altered versions of matrix $A$, correspondingly, while ||.||||.||| | . | | signifies the Euclidean distance among these cross-attention maps. Furthermore, the inclusion of the regularization term $\gamma\mathcal{L}_{norm}$ derived from self-guidance (Epstein et al. 2023), strives to prevent overfitting, where $\gamma$ denotes a weighting parameter.
Figure 4: The results of comparison. Each column showcases six prompts across three datasets, the final column depicting the saliency map of the result image generated from the mushroom prompt. The red-dotted area denotes the planned blank region. Note that some methods fail to follow the orange-highlighted words in the prompt, leading to semantic loss.
Dataset Metrics Stable Diffusion Dall-E 3 AnyText Desigen Ours
P2P Template Saliency IOU $\downarrow$ 29.89 52.64 54.34 30.62 28.37
TV Loss $\downarrow$ 14.11 18.02 22.55 13.70 13.29
VTCM $\uparrow$ 2.95 1.92 1.75 2.92 3.09
DiffuisonDB Saliency IOU $\downarrow$ 30.11 56.82 55.88 30.30 27.80
TV Loss $\downarrow$ 12.30 21.16 20.39 11.99 11.55
VTCM $\uparrow$ 3.19 1.67 1.79 3.20 3.37
Synthetics Prompt Saliency IOU $\downarrow$ 31.42 51.52 53.24 31.57 30.83
TV Loss $\downarrow$ 15.61 17.85 21.46 15.63 15.17
VTCM $\uparrow$ 2.73 1.98 1.79 2.67 2.77
Table 1: Quantitative comparison of our method with other methods (excluding CLIP Score). Bold indicates the best scores.
 Dataset Dall-E 3 AnyText Desigen Ours
P2P Template 2.53 0.32 0.55 0.30
DiffuisonDB 2.04 1.04 0.34 0.31
Synthetics Prompt 2.25 1.02 0.51 0.29
Table 2: Quantitative comparison of CLIP Scores Loss for different methods. Bold indicates the best scores.
Experiments
The evaluation is structured into quantitative and qualitative analyses, alongside an ablation study to understand the contribution of individual components of our model.
Implementation Details.
@@
Dataset for Evaluation.
Our evaluation contains 27,000 images generated from 2,700 unique prompts, each tested in ten different random region $R$. The dataset combined synthesized prompts generated by ChatGPT, and 700 prompts from the Prompt2Prompt template (Hertz et al. 2022) designed for attention guidance, focusing on specific objects and their spatial relationships. Additionally, we included 1,000 DiffusionDB prompts (Wang et al. 2022), chosen for their real-world complexity. This diverse and comprehensive dataset, spanning synthetic to user-generated prompts, provided a broad test ground to evaluate the efficacy of our TextCenGen method in various T2I scenarios.
Comparison with Existing Methods
We compared TextCenGen with several potential models to evaluate its efficiency. The baseline models included: Native Stable Diffusion (Rombach et al. 2022), Dall-E 3 (Ramesh et al. 2022), AnyText (Tuo et al. 2023) and Desigen (Weng et al. 2024). Dall-E used the prompt ”text-friendly in the position” to specify the region $R$. Similar to AnyText, we chose to randomly generate several masks in a fixed pattern across the map to simulate regions need to be edited. More detail can be found in appendix.